# Quantum coherence and the invisible Universe: Subradiance as a dark matter mechanism

### v2

@@ 2.1 Entangled quantum states and coherent behavior
It is shown in Appendix A.2 that for $n\gg 1$ the energy trapping efficiency is $\eta\approx 1-\sqrt{2/n}$. It is therefore clear that the combination of coherence and the entanglement of a large number of atoms can potentially lead to a significant amount of energy trapping within the gas. But one of our goals is to assess the level of energy trapping within the gas for finite and more realistic atomic separations in the interstellar medium. It is already obvious through our study of the two-atom system that the general case where $kz_{0}\neq 0$ will lead to a non-zero transition rate from the |$\Ket{0,0}_{\theta}$ ‘dark’ state (see equations 4 and 32). This has for implication that all the energy initially stored in the system will eventually leak away. That is, $\eta=0$ in the steady-state and no energy trapping is realized in the long run.
@@ 3 Discussion
Another important property of our darkened atomic hydrogen gas is discussed in Appendix D. There, we calculate the cross-section for collisions between two two-atom systems, both being entangled as discussed in Sec. 2.1. Interestingly, it is found that this collision cross-section vanishes because of destructive interference due to the entangled nature of the states the systems find themselves in after the transient regime. Although this analysis needs to be generalized, this result is consistent with astronomical observations of the so-called bullet cluster and similar objects [32]. These observations put a constraint on the ratio of the collision cross-section to the mass of a purported dark matter (point) particle to $\sigma/M\lesssim 1\,\mathrm{cm^{2}\,g^{-1}}$ [27]. Although the corresponding figure for atomic hydrogen is $\sim 10^{9}\,\mathrm{cm^{2}\,g^{-1}}$, it does not apply in our case. Our analysis does not suggest that atomic hydrogen is a dark matter particle, but rather that the cooperative behavior of a large number of entangled hydrogen atoms is consistent with what is expected for dark matter. While individual atoms may still collide, entangled systems of atoms will be essentially collisionless.
@@ 4 Methods
For the atomic hydrogen slab of density $1\,\mathrm{cm^{-3}}$ and temperature $T_{\mathrm{kin}}=10$ K considered in Secs. 2 and 2.3.2, thermal motions bring an average time-scale between collisions of about $2.5\times 10^{9}$ s, which we ascribe to $T_{2}$. Since we are considering radiation from the 21 cm line, which has an excitation temperature $T_{\mathrm{21\,cm}}\simeq 0.068\,\mathrm{K}\ll 10$ K, we set, for simplicity, $T_{1}\approx T_{2}=2.5\times 10^{9}$ s (i.e., we assume that inelastic and elastic collisions happen at similar rates and are the sole sources of relaxation and dephasing in the gas, respectively).
@@ Acknowledgements
Martin Houde1,∗ and Fereshteh Rajabi2
1,∗Department of Physics and Astronomy, The University of Western Ontario,
@@ A.2 Energy trapping efficiency for $n$ atoms
For the two-atom problem only two permutation symmetries are available for the Dicke states: the totally symmetric species is shared by the $r=1$ triplet (equations 1-3), while the lone $r=0$ singlet is anti-symmetric (equation 4). In the more general case where the gas is composed of an arbitrary number of atoms $n$ a larger number of symmetries, tied to the cooperative number $r$, exist for the Dicke states.
@@
One can readily verify that $\eta=1/2$ for the two-atom case, while it is $0.69$ for $n=10$, $0.88$ for $n=100$, etc. Furthermore, it is known that $\overline{r\left(r+1\right)}\simeq m^{2}+n/2$ whenever $n\gg 1$ and $\hbar\omega_{0}\ll k_{\mathrm{B}}T_{\mathrm{kin}}$ we have [1], which from equation (42) yields $\eta\approx 1-\sqrt{2/n}$ when $m=0$.
@@ Appendix D Collision cross-section
|$\Ket{\psi_{+}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}-\hbar\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}+e^{-i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (85)
@@
|$\Ket{\psi_{-}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}-\hbar\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}-e^{-i\hbar\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (86)
@@
If two atoms enter into an elastic collision, say the first atoms of both pairs, then under the Born approximation each term will bring a scattering amplitude [39]
@@
with $\mathbf{k}_{\mathrm{in}}=\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)/\hbar$, $\mathbf{k}_{\mathrm{out}}$ the scattered wave vector and $m_{\mathrm{r}}$ the reduced mass. It therefore follows from equation (87) that the collision cross-section $\sigma\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=0$ since it consists of the square of the sum of the four corresponding scattering amplitudes.
Although such calculations need to be generalized, we find that two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.

### v3

@@ 2.1 Entangled quantum states and coherent behavior
It is shown in Appendix A.2 that for $n\gg 1$ the energy trapping efficiency is $\eta\approx 1-\sqrt{2/n}$ when starting from $m=0$. It is therefore clear that the combination of coherence and the entanglement of a large number of atoms can potentially lead to a significant amount of energy trapping within the gas. But one of our goals is to assess the level of energy trapping within the gas for finite and more realistic atomic separations in the interstellar medium. It is already obvious through our study of the two-atom system that the general case where $kz_{0}\neq 0$ will lead to a non-zero transition rate from the |$\Ket{0,0}_{\theta}$ ‘dark’ state (see equations 4 and 32). This has for implication that all the energy initially stored in the system will eventually leak away. That is, $\eta=0$ in the steady-state and no energy trapping is realized in the long run.
@@ 3 Discussion
Another important property of our darkened atomic hydrogen gas is discussed in Appendix D. There, we calculate the cross-section for high-speed collisions between two two-atom systems (i.e., when the kinetic energy is far greater than the transition excitation energy $\hbar\omega_{0}$), both being entangled as discussed in Sec. 2.1. Interestingly, it is found that in such circumstances the collision cross-section will likely vanish because of destructive interference due to the entangled nature of the states the systems find themselves in after the transient regime. Although this analysis needs to be generalized, this result is consistent with astronomical observations of the so-called bullet cluster (with collision speeds on the order of $10^{3}\,\mathrm{km\,s^{-1}}$) and similar objects [32]. These observations put a constraint on the ratio of the collision cross-section to the mass of a purported dark matter (point) particle to $\sigma/M\lesssim 1\,\mathrm{cm^{2}\,g^{-1}}$ [27]. Although the corresponding figure for atomic hydrogen is $\sim 10^{9}\,\mathrm{cm^{2}\,g^{-1}}$, it does not apply in our case. Our analysis does not suggest that atomic hydrogen is a dark matter particle, but rather that the cooperative behavior of a large number of entangled hydrogen atoms is consistent with what is expected for dark matter. While individual atoms may still collide, entangled systems of atoms will be essentially collisionless.
@@ 4 Methods
For the atomic hydrogen slab of density $1\,\mathrm{cm^{-3}}$ and temperature $T_{\mathrm{kin}}=10$ K considered in Secs. 2.3.1 and 2.3.2, thermal motions bring an average time-scale between collisions of about $2.5\times 10^{9}$ s, which we ascribe to $T_{2}$. Since we are considering radiation from the 21 cm line, which has an excitation temperature $T_{\mathrm{21\,cm}}\simeq 0.068\,\mathrm{K}\ll 10$ K, we set, for simplicity, $T_{1}\approx T_{2}=2.5\times 10^{9}$ s (i.e., we assume that inelastic and elastic collisions happen at similar rates and are the sole sources of relaxation and dephasing in the gas, respectively).
@@ Acknowledgements
[…]
${}^{1^{\ast}}$epartment of Physics and Astronomy, The University of Western Ontario,
@@ A.2 Energy trapping efficiency for $n$ atoms
For the two-atom problem only two permutation symmetries are available for the Dicke states: the totally symmetric species is shared by the $r=1$ triplet (equations 1-3), while the lone $r=0$ singlet is anti-symmetric (equation 4). In the more general case where the gas is composed of an arbitrary number of atoms $n$ a larger number of symmetries, tied to the cooperative number $r$, exist for the Dicke states (see Fig. 1).
@@
One can readily verify that $\eta=1/2$ for the two-atom case, while it is $0.69$ for $n=10$, $0.88$ for $n=100$, etc. Furthermore, it is known that $\overline{r\left(r+1\right)}\simeq m^{2}+n/2$ whenever $n\gg 1$ and $\hbar\omega_{0}\ll k_{\mathrm{B}}T_{\mathrm{kin}}$ [1], which from equation (42) yields $\eta\approx 1-\sqrt{2/n}$ when $m=0$.
@@ Appendix D Collision cross-section
|$\Ket{\psi_{+}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}/\hbar-\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}+e^{-i\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (85)
@@
|$\Ket{\psi_{-}}$ $\propto\int d^{3}R\,e^{i\left(\mathbf{P}/\hbar-\mathbf{k}_{r}\right)\cdot\mathbf{R}}\,\Ket{\mathbf{R}}\int d^{3}\Delta r\,\Ket{\Delta\mathbf{r}}$
×$\times\frac{1}{\sqrt{2}}\left(e^{i\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ab}-e^{-i\mathbf{k}_{r}\cdot\Delta\mathbf{r}/2}\Ket{ba}\right)\Ket{0}.$ (86)
@@
If two atoms enter into an elastic collision, say the first atoms of both pairs, then under the Born approximation each term will bring a scattering amplitude of the type [38]
@@
with $\mathbf{k}_{\mathrm{in}}=\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)/\hbar$, $\mathbf{k}_{\mathrm{out}}$ the scattered wave vector and $m_{\mathrm{r}}$ the reduced mass. For high relative speed between the colliding atoms, i.e., when $\left|\mathbf{P}_{B}-\mathbf{P}_{A}\right|^{2}/2m_{\mathrm{r}}\gg\hbar\omega_{0}$, it is expected that the scattering amplitude will be independent of their internal states. It therefore follows from equation (87) that under these circumstances the collision cross-section $\sigma\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=0$ since it consists of the square of the sum of the four corresponding scattering amplitudes.
Although such calculations need to be generalized, we find that at high relative speed two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.